OpenStudy (anonymous):

20 sin [2π(x-20)] +20 = 10. How do I find the solutions for x?

OpenStudy (anonymous):

\[20\sin \left\{ 2\pi \left( x-20 \right) \right\}=10-20\] \[\sin \left\{ 2\pi \left( x-20 \right) \right\}=\frac{ -10 }{20}=\frac{ -1 }{ 2 }=-\sin \frac{ \pi}{6 }\] \[=\sin \left( \pi+\frac{ \pi }{6 } \right),\sin \left( 2\pi-\frac{ \pi }{ 6 } \right)\] \[=\sin \frac{ 7\pi }{ 6},\sin \frac{ 13\pi }{ 6 }\] \[2\pi \left( x-20 \right)=\frac{ 7\pi }{6 },\frac{ 13\pi }{6 }\] \[x-20=\frac{ 7 }{ 12 },\frac{ 13 }{12 }\] calculate x

OpenStudy (anonymous):

omfg thank you